OpenStudy (anonymous):

ANSWER IN 5 MINUTES GETS A MEDAL + FAN! Which is an equation in point-slope form for the given point and slope? Point: 8, -4 ; Slope: -5/6 @iGreen @kirbykirby @pgpilot326 @paki

OpenStudy (anonymous):

Do you know the formula of point-slope form? \[y-y1=m(x-x1)\] Simply replace the values by the number.

OpenStudy (anonymous):

@im.celibate

OpenStudy (anonymous):

I forgot.. but here arre the choices(: http://static.k12.com/bank_packages/files/media/mathml_1dcb67f9683450ff9c7236824cdf92e30c8d8e01_1.gif

OpenStudy (anonymous):

That's the formula that you applied, correct?

OpenStudy (anonymous):

never mind.. I got it could you help me with another one?

OpenStudy (anonymous):

\[y+4=-\frac{ 5 }{ 6 }(x-8)\]

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Thanks, What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3?

OpenStudy (anonymous):

@vzfreakz

OpenStudy (anonymous):

Do the same process like you did on the first one. Replace the values by the numbers.

OpenStudy (anonymous):

i don'tundrestand

OpenStudy (anonymous):

@vzfreakz

OpenStudy (anonymous):

You shown me your equation, and you simply replace the values (x and y), am I mistaking? That same process that is the same on your next question. \[y-4=-3(x-1)\]